[Hallc_running] [New Logentry] Hall C RC update: 10/20/2021

gaskelld at jlab.org gaskelld at jlab.org
Wed Oct 20 08:25:02 EDT 2021

Logentry Text:
Day shift yesterday was very productive after somewhat unstable beam on the shift before.  Unfortunately, an issue with the LCW in the South Access Building took down the accelerator early swing shift.  After repair of the LCW system, the conductivity of the water remained too high until this morning, so there was no beam for most of swing and all of owl.

Today is a planned accelerator maintenance day.  Given the planned work in the tunnel, RF recovery, and beam restoration time, beam is not expected to the halls before midnight.

Hall C is now in Controlled Access - please let the RC or the Hall C shift leader know before you enter the hall.

Given that beam is not expected for day and swing shifts, Shift Leaders are on-call today (please make sure the target operator can contact you if needed).  Shifts should be fully staffed again starting Thursday owl.

Once beam returns, please follow the abbreviated beam recovery procedure described here:

and then continue production running on LH2 at 42 uA.

There will be an RC meeting this morning at 10 am:
Bluejeans link: https://bluejeans.com/861439920
RC meeting page: https://hallcweb.jlab.org/wiki/index.php/RC_Meeting,_Wednesday,_October_20_,_2021

Finally, Dave Mack will take over as RC today. 


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3928536
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211020/7f636b38/attachment.html>

More information about the Hallc_running mailing list